OpenStudy (anonymous):

a ball of mass 10g is hung from the end of light vertical spring and set oscillating with an amplitude of 10cm. the period of the oscillation is 1s. if the ball is now replaced by another with a mass of 40g and again set oscillating with the same amplitude,what will be its new period of oscillation?

OpenStudy (anonymous):

i think this \[\omega _{1}=(\sqrt{k/m _{1}})\rightarrow k=m _{1}\omega _{1}^{2}\] \[T _{2}=2\pi \sqrt{(m _{2}/k)}\]\[T _{2}=2\pi \sqrt{(m _{2}/m _{1}\omega _{1}^{2})}\] i reached to\[T _{2}=2\]